The balance sheet for Stud Clothiers is shown next. Sales for the year were $3,490,000, with 75 percent of sales sold on credit. STUD CLOTHIERS Balance Sheet 20X1 Assets Liabilities and Equity $262,000 148,000 178,000 100,000 $ 38,000 Accounts payable Accounts receivable Inventory Plant and equipment 292,000 248,000 500,000 Accrued taxes Bonds payable (long-term) Common stock Paid-in capital Retained earnings 150,000 240,000 $ 1,078,000 $ 1,078,000 Total assets Total liabilities and equity Compute the following ratios: (Use a 360-day year. Do not round intermediate calculations. Round your answers to 2 decimal places. Input your debt-to-total assets answer as a percent rounded to 2 decimal places.) a. Current ratio b. Quick ratio c. Debt-to-total-assets ratio d Asset turnover e. Average collection period times times times days
Answer #1
a. Current ratio           1.41 times
Working:
# 1 Current assets = Cash +Accounts receivable + Inventory
= 38000+292000+248000
= $          5,78,000
# 2 Current Liabilities = Accounts Payable + Accrued Taxes
= 262000+148000
= $          4,10,000
# 3 Current ratio = Current Assets / Current Liabilities
= $          5,78,000 / $          4,10,000
=                      1.41
b. Quick ratio           0.80 times
Working:
Quick assets = Cash + Accounts Receivable
= 38000+292000
= $          3,30,000
Quick Ratio = Quick assets / Current Liabilities
= $          3,30,000 / $          4,10,000
=                      0.80
c. Debt to total assets ratio 54.55%
Working:
# 1 Total Debt = Current Liabilities + Bonds Payable
= $          4,10,000 + $          1,78,000
= $          5,88,000
# 2 Debt to total assets ratio = Total debt / Total assets
= $          5,88,000 / $       10,78,000
= 54.55%
d. Asset Turnover           3.24 times
Working:
Asset Turnover = Net Sales / Total assets
= $       34,90,000 / $       10,78,000
=                      3.24
e. Average collection period         40.16 Days
Working:
# 1 Credit Sales = 3490000*75%
= $       26,17,500
# 2 Accounts receivable turnover = Net credit sales /Accounts receivable
= $       26,17,500 / $          2,92,000
=                      8.96
# 3 Average collection period = Days in a year / Accounts Receivable turnover ratio
= 360 /                      8.96
=                    40.16